OpenStudy (bimbim2697):

Find the water pressure (in kPa) at the 25.0-m level of a water tower containing water 50.0 m deep.

OpenStudy (irishboy123):

\(P = P_{atm} + \rho_w g h\), a sub-set of the Bernoulli Equation \(\rho_w\) is the density of water In this case: \(P = 101325 + 1000 \cdot 9.81 \cdot 25 = 346,575\) Units in Pascals !! for kPa you need to convert to get the answer